Can someone please help me with this question? I tried doing it and I got 2.4x10^8, but I am not sure if that is right. Thanks in advanced.

A 1/10 serial dilution of a bacterial culture was performed in tubes "A" through "E" (dilutions were made in that order beginning with tube "A" like in the handout slide titled "Spread Plating"). 100 microliters of the diluted culture from the final tube "E" was spread onto a plate. Following incubation of the plate, growth of the bacteria produced 240 colonies on that plate. What is the concentration of viable cells in tube "C" (in cells per milliliter)?
Answer #1

Dilution factor plate E= 1/10 x 1/10 x 1/10 x 1/10 x 1/10

=10-5

Number of colonies= 240

Volume of sample =0.1 mL to the plate

cfu/ml = (no. of colonies)/( dilution factor x volume of culture plate)

cfu/ml = (240)/ (10-5 x 0.1)

cfu/ml = 2.4 x 108 cfu/mL

the concentration of viable cells in tube "B" = Dilution factor (tube B) x original number of cells

=10-2 x 2.4 x 108 cfu/mL

=2.4 x 106 cells/mL